# Character tables are ideal Perron similarities

### v1

@@ 1 Introduction
is called the (Perron) spectracone of $S$, and the polytope
@@ 2 Background & Notation
If $X\subseteq\mathbb{F}^{n}$, then the conical hull of $X$, denoted by $\operatorname{\mathsf{coni}}X=\operatorname{\mathsf{coni}}(X)$, is defined by
$\operatorname{\mathsf{coni}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,~\alpha_{k}\geqslant 0\right\},&X\neq\varnothing.\end{cases}$
i.e., when $X$ is nonempty, $\operatorname{\mathsf{coni}}X$ consists of all conical combinations. Similarly, the convex hull of $X$, denoted by $\operatorname{\mathsf{conv}}{X}=\operatorname{\mathsf{conv}}(X)$ is defined by
$\operatorname{\mathsf{conv}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,\ \sum_{k=1}^{m}\alpha_{k}=1,\ \alpha_{k}\geqslant 0\right\},&X\neq\varnothing\end{cases}.$
@@ 2.1 Characters & Representations
[…]
charlinearcombo If $\rho_{1},\ldots,\rho_{n}$ are the distinct irreducible representations of $G$ and $\chi_{1},\ldots,\chi_{n}$ are their respective characters, then
@@ 3 Spectral Polyhedra & Perron Similarities
[…]
spectratope If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then:
@@
Reall from the introduction that an invertible matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$. It is known that $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $r_{i}\in\mathcal{C}(S),\ \forall i\in\left[n\right]$, i.e., every row, viewed as a column vector, belongs to its spectracone [11, Theorem 4.21].
@@
[…]
rhc If $S\in\mathsf{M}_{m\times n}({\mathbb{C}})$, then $S$ is called row Hadamard conic (RHC) if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall i,j\in\left[m\right]$.
@@
[…]
simple If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $x^{\top}S^{-1}\geqslant 0$ if and only if $x\in\mathcal{C}_{r}(S)$.
@@
[…]
rhcrows If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $r_{i}\in\mathcal{C}(S)$ if and only if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),\ \forall j\in\left[n\right]$.
@@
$\displaystyle\Longleftrightarrow r_{j}\circ r_{i}=r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall j\in\left[n$
@@
[…]
@@
[…]
RHC ideal If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $S$ is RHC.
@@
[…]
epairconstruct Supoose that $S\in\mathsf{GL}_{n}(\mathbb{C})$, $\lambda\in\mathbb{C}^{n}$, and $A\coloneqq SD_{\lambda}S^{-1}$. If $\exists k\in[n]$ such that $v\coloneqq Se_{k}$ is totally nonzero, then $(\lambda_{k},e)$ is an eigenpair of $D_{v^{-1}}AD_{v}$.
@@
[…]
P(S)contain If $S\in\mathsf{GL}_{n}(\mathbb{C})$ is a Perron similarity, then there exists $k\in[n]$ such that $\mathcal{P}(S)\subseteq\mathcal{P}(D_{v^{-1}}S)$, where $v\coloneqq Se_{k}$.
@@
[…]
@@
[…]
@@ 4 Main Results
$\displaystyle=\frac{q_{i,1}\overline{q_{j,1}}}{|C_{G}(g_{1})|}$
$\displaystyle=\frac{\dim(\rho_{i})\dim(\rho_{j})}{|C_{G}(g_{1})|}>0.$
@@
extremerayQ The matrix $Q$ is ideal.
Proof.
Since
[…]
it follows that $e_{1}^{\top}Q=e^{\top}$, i.e., $e\in\mathcal{C}_{r}(Q)$.
Suppose that $i,j\in[n]$. For ease of notation, let
$\alpha_{k}\coloneqq\langle\chi_{\rho_{i}\otimes\rho_{j}},\chi_{k}\rangle,\ \forall k\in\left[n\right]$
and $\alpha\coloneqq\begin{bmatrix}\alpha_{1}&\cdots&\alpha_{n}\end{bmatrix}^{\top}$. With the above in mind,
$\displaystyle r_{i}^{\top}\circ r_{j}^{\top}$ $\displaystyle=\begin{bmatrix}\chi_{i}(g_{1})\chi_{j}(g_{1})&\cdots&\chi_{i}(g_{n})\chi_{j}(g_{n})\end{bmatrix}$
$\displaystyle=\begin{bmatrix}\chi_{\rho_{i}\otimes\rho_{j}}(g_{1})&\cdots&\chi_{\rho_{i}\otimes\rho_{j}}(g_{n})\end{bmatrix}$ [Eq. 2.2]
$\displaystyle=\begin{bmatrix}\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{1})&\cdots&\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{n})\end{bmatrix}$ [Eq. 2.5]
$\displaystyle=\alpha^{\top}Q.$
[…]
@@
[…]
ineqredux If $x\in\mathbb{C}^{n}$, then $M_{x}(Q)\geqslant 0$ if and only if
@@
To demonstrate the necessity of Condition 4, notice that if $M_{x}(Q)\geqslant 0$ and $i\in\left[n\right]$, then
@@
Conversely, if Condition 4 holds and $i,j\in[n]$, then
@@
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\chi_{\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.3]
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\rho_{i}\otimes\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.2]
[…]
@@
$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)$
[…]
which establishes the sufficiency of Condition 4. ∎
@@
$\displaystyle=Q(\overline{D^{-1}}D_{\overline{x}}D_{x}D^{-1})Q^{\ast}$
$\displaystyle=Q(D_{x}D^{-1}\overline{D^{-1}}D_{\overline{x}})Q^{\ast}$
@@ 5 Volume of the Projected Spectratope of a Real Character Table
[…]
scalingConeLemma If $S$ is ideal and $v>0$, then $D_{v}S$ is ideal.
@@
[…]
@@
[…]
volumeSpectratope The volume of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
@@ 6 Examples
[…]
@@
[…]
@@
$\displaystyle V$ $\displaystyle=\frac{\sqrt{4^{4}}}{(3-1)!\cdot 1^{4}}=\frac{8}{3}.$

### v2

@@ 1 Introduction
where $D_{x}$ is the diagonal matrix whose $(k,k)$-entry is $x_{k}$, is called the (Perron) spectracone of $S$, and the polytope
@@ 2 Background & Notation
If $\varnothing\subset X\subseteq\mathbb{F}^{n}$, then the conical hull of $X$, denoted by $\operatorname{\mathsf{coni}}X=\operatorname{\mathsf{coni}}(X)$, is defined by
$\operatorname{\mathsf{coni}}{X}=\left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,~\alpha_{k}\geqslant 0\right\}$
i.e., $\operatorname{\mathsf{coni}}X$ consists of all conical combinations. Similarly, the convex hull of $X$, denoted by $\operatorname{\mathsf{conv}}{X}=\operatorname{\mathsf{conv}}(X)$ is defined by
$\operatorname{\mathsf{conv}}{X}=\left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,\ \sum_{k=1}^{m}\alpha_{k}=1,\ \alpha_{k}\geqslant 0\right\}.$
@@ 2.1 Characters & Representations
If $\rho_{1},\ldots,\rho_{n}$ are the distinct irreducible representations of $G$ and $\chi_{1},\ldots,\chi_{n}$ are their respective characters, then
@@ 3 Spectral Polyhedra & Perron Similarities
If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then:
@@
Recall from the introduction that an invertible matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$. It is known that $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $r_{i}\in\mathcal{C}(S),\ \forall i\in\left[n\right]$, i.e., every row, viewed as a column vector, belongs to its spectracone [11, Theorem 4.21].
@@
If $S\in\mathsf{M}_{m\times n}({\mathbb{C}})$, then $S$ is called row Hadamard conic (RHC) if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall i,j\in\left[m\right]$.
@@
If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $x^{\top}S^{-1}\geqslant 0$ if and only if $x\in\mathcal{C}_{r}(S)$.
@@
If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $r_{i}\in\mathcal{C}(S)$ if and only if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),\ \forall j\in\left[n\right]$.
@@
$\displaystyle\Longleftrightarrow r_{j}\circ r_{i}=r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall j\in\left[n\right]$ [?THM? LABEL:simple]
@@
The following characterization is an immediate result of ?THM? LABEL:rhcrows.
@@
If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $S$ is RHC.
@@
Supoose that $S\in\mathsf{GL}_{n}(\mathbb{C})$, $\lambda\in\mathbb{C}^{n}$, and $A\coloneqq SD_{\lambda}S^{-1}$. If $\exists k\in[n]$ such that $v\coloneqq Se_{k}$ is totally nonzero, then $(\lambda_{k},e)$ is an eigenpair of $D_{v^{-1}}AD_{v}$.
@@
If $S\in\mathsf{GL}_{n}(\mathbb{C})$ is a Perron similarity, then there exists $k\in[n]$ such that $\mathcal{P}(S)\subseteq\mathcal{P}(D_{v^{-1}}S)$, where $v\coloneqq Se_{k}$.
@@
If $z\in\mathcal{P}(S)$, then $M_{z}(S)$ is a stochastic matrix. Since $(z_{k},v)$ is an eigenpair, it follows that $(z_{k},x)$ is an eigenpair. As $M_{z}(S)$ is a nonnegative matrix having a positive eigenvector, it is necessarily the case that $z_{k}=\rho(M_{z}(S))$ [6, Corollary 8.1.30]. As $M_{z}(S)$ is stochastic, $z_{k}=1$. Finally, ?THM? LABEL:epairconstruct ensures that $(z_{k},e)=(1,e)$ is an eigenpair of $M_{z}(D_{v^{-1}}S)$. ∎
@@
The inclusion in ?THM? LABEL:P(S)contain can be strict: if $S=\begin{bmatrix}1&1\\ 2&-2\end{bmatrix}$, then the matrix
@@ 4 Main Results
$\displaystyle=\frac{q_{i,1}\overline{q_{j,1}}}{|C_{G}(e_{G})|}$
$\displaystyle=\frac{\dim(\rho_{i})\dim(\rho_{j})}{|G|}>0.$
@@
The result is clear because the $k$th entry of $r_{i}^{\top}\circ r_{j}^{\top}$ is
$\chi_{i}(g_{k})\cdot\chi_{j}(g_{k})=\chi_{\rho_{i}\otimes\rho_{j}}(g_{k})=\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}},\chi_{\ell}\rangle\chi_{\ell}(g_{k}),$
which demonstrates that $Q$ is RHC and thus ideal by ?THM? LABEL:RHC_ideal. ∎
Whereas ?THM? LABEL:extremerayQ yields a description of $\mathcal{C}(Q)$ as the conical hull of a set of linearly independent vectors, the subsequent result describes $\mathcal{C}(Q)$ in terms of linear inequalities and demonstrates that not all of the $n^{2}$ inequalities of the matrix equation $M_{x}(Q)\geqslant 0$ are required to specify the cone. The linear inequalities are novel sufficient conditions for realizability.
@@
If $x\in\mathbb{C}^{n}$, then $M_{x}(Q)\geqslant 0$ if and only if
@@
To demonstrate the necessity of Condition 4.2, notice that if $M_{x}(Q)\geqslant 0$ and $i\in\left[n\right]$, then
@@
Conversely, if Condition 4.2 holds and $i,j\in[n]$, then following properties of representations and ?THM? LABEL:charlinearcombo
@@
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\rho_{i}\otimes\rho_{j}^{*}}(g_{k})x_{k}$
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}\left(\sum_{\ell=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\chi_{\ell}(g_{k})x_{k}\right)$
@@
$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)\geqslant 0,$
which establishes the sufficiency of Condition 4.2. ∎
@@
$\displaystyle=Q(\overline{D^{-1}}D_{\overline{x}}DD_{x}D^{-1})Q^{\ast}$
$\displaystyle=Q(D_{x}D^{-1}D\overline{D^{-1}}D_{\overline{x}})Q^{\ast}$
@@ 5 Volume of the Projected Spectratope of a Real Character Table
If $S$ is ideal and $v>0$, then $D_{v}S$ is ideal.
@@
If $v\coloneqq Qe_{1}$, then, since $Q$ is ideal, it follows that $D_{v^{-1}}Q$ is ideal and $(D_{v^{-1}}Q)e_{1}=e$. Recall, by ?THM? LABEL:P(S)contain, that $\mathcal{P}(Q)\subseteq\mathcal{P}(D_{v^{-1}}Q)$.
@@
The volume of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
@@ 6 Examples
By ?THM? LABEL:ineqredux, the half-spaces that determine the polyhedral cone $\mathcal{C}(Q)$ are $x_{1}+3x_{2}+2x_{3}\geqslant 0$, $x_{1}-3x_{2}+2x_{3}\geqslant 0$, and $2(x_{1}-x_{3})\geqslant 0$. Indeed, the inequality $2(2x_{1}+x_{3})\geqslant 0$ is redundant since
@@
If $x\in\mathbb{R}^{4}$, then, following ?THM? LABEL:ineqredux, $M_{x}(H_{4})\geqslant 0$ if and only if
@@
$\displaystyle V$ $\displaystyle=\frac{\sqrt{4^{4}}}{(4-1)!\cdot 1^{4}}=\frac{8}{3}.$